# A Canonical Quantization of Poisson Manifolds: a 2-Groupoid Scheme

### v1

@@ Abstract
We canonically quantize a Poisson manifold to a Lie 2-groupoid, complete with a quantization map, and show that it relates geometric and deformation quantization: the perturbative expansion in of the (formal) convolution of two quantized functions yields Kontsevich’s star product. Meanwhile, we can push forward this quantization map (by integrating over homotopies of paths) to obtain a quantization map in traditional geometric quantization. This gives a polarization-free, path integral definition of the quantization map, which does not have a prior definition for Poisson manifolds, and which only has a partial prescription for symplectic manifolds. We construct conventional quantum mechanics from this perspective.
@@ 0 Introduction
Kontsevich proved that every Poisson manifold admits a formal deformation quantization, obtained using a sequence of bidifferential operators, called a star product (see [14]). That is, there is an associative product on which deforms the product on in such a way that
@@
A non-perturbative approach to quantization, which was initiated by Weinstein in [25], is geometric quantization of Poisson manifolds. This quantization method aims to define a -algebra of observables — it is given by polarized sections of a twisted convolution algebra of the symplectic groupoid (the groupoid which integrates the Poisson structure, [10]). In the symplectic case, this -algebra should act on polarized sections of the prequantum line bundle.222For a discussion on when a star product acts on the space of sections of such a quantization, see [17], [18].
@@
Geometric quantization is much less understood than formal deformation quantization (which is consistent with the current understanding of QFT) and is incomplete as it doesn’t prescribe a quantization map.333Here, we mean an injective ∗-linear map from to the -algebra of observables that determines a star product. In this sense, the difficulty of geometric quantization is complementary to the difficulty of formal deformation quantization: one describes a -algebra but not a quantization map, the other describes a formal quantization map but not a -algebra.
@@
In some examples (eg. [1],[3],[21]) it has been shown that there exists a quantization map which leads to a formal deformation quantization, but an explanation of why and when this occurs is lacking. Even in traditional geometric quantization of symplectic manifolds the quantization map is, at best, defined only on a small subspace (and assumes a polarization exists). A goal is to remove some of the mystery from the theory.
@@
The constructions we make use results based on the Poisson sigma model in [2], [5]. Formally, we define a quantization map by using the Lie 2-groupoid integrating a Poisson manifold ([26]), which involves computing a path integral over homotopies of algebroid paths. Noncommutativty for is due distinct points in phase space being isomorphic in the 2-groupoid.
Our main example is the pointed Kähler manifold where the quantization of a smooth function is given by
@@
where is the (signed) area of the polygon determined by and the integration is over (eq. 4.0.8).
The canonical connection on the prequantum line bundle determines a representation of these quantized functions as operators on and polarized sections form subrepresentations, for all covariantly constant Lagrangian polarizations. In particular, we get the position, momentum and Segal-Bargmann (holomorphic) representations.444In fact, it is nicer to view these irreducible representations as embedded inside this larger one. With the Kostant–Souriau prescription, only a small subspace of functions prequantize to operators which preserve polarized sections, whereas all of our operators preserve them (at least in this example).
@@
We view the role of Lagrangian polarizations in the usual definitions of geometric quantization as a tool to help guess a quantization map, but not necessarily as an essential ingredient. This is consistent with Kontsevich’s and Fedosov’s ([8]) constructions of the star product, where polarizations are absent. We will only use them to construct irreducible representations.
@@ 0.1 Lie 2-Groupoid Approach to Quantization and Main Results
Its space of sections, denoted is formally a -algebra — its product is given by convolution (this involves a path integral) and its involution ∗ is complex conjugation. It’s norm is (formally) the -norm of the left regular representation.555Every algebra acts on itself on the left, forming a representation called the left regular representation. There is well-defined, canonical quantization map
@@
(formal) Let be a Poisson maniold such that the convolution of two sections of is smooth.666This should be true in a wide class of examples, including regular Poisson manifolds satisfying an integrality condition. We have that
@@
(0.1.4) of the Poisson sigma model on the disk ([5]). Therefore, asymptotically at
@@
0.2 *Non-Perturbative Tangential Deformation Quantizations
Here, we will discuss non-perturbative deformation quantizations. The definition we use is essentially the most general definition that is compatible with both -algebraic quantization and formal deformation quantization. It sits between Rieffel’s strict quantization ([19]) and his strict deformation quantization (one may come back to this section later).
@@
be injective and ∗- linear such that its image generates Furthermore, assume that is the inclusion map.
We say that is a (non-perturbative) deformation quantization of if there is a star product on such that, for all
@@
Every regular Poisson manifold has a tangential, formal deformation quantization,777A star product is tangential if the sequence of bidifferential operators are tangent to the symplectic leaves. see the work of Fedosov [8], also see [12], [24]. (there are a number of examples of non-regular Poisson manifolds which also have such a quantization, as discussed in [16]).888A quantization which isn’t tangential exhibits a kind of tunneling. The maps discussed in section 0.1 have a tangential property since the orbits of the Lie 2-groupoid are the symplectic leaves.
@@
where the supremum is over all symplectic leaves and is the -norm of pulled back to 999With respect to the measure induced by the symplectic form. We denote this space by
@@
This would give a a non-perturbative form of Kontsevich’s star product on such Poisson manifolds.101010For examples of non-tangential quantizations, see the quantizations of the dual of nilpotent Lie algebras [20].
@@
In formal deformation quantization, it’s really the gauge equivalence class of the formal quantization map that’s important. Similarly, we can define two quantization maps to be gauge equivalent if there is an automorphism of -algebras intertwining them.
@@ 1 The Star Product and the Poisson Sigma Model
The star product is then given by111111More precisely, we can say that the asymptotic expansions at is Kontsevich’s star product.
@@ 1.1 The Tangential Star Product
A lattice approach to defining these path integrals is described in [16], see appendix 5. On a Kähler manifold this kernel should be canonically defined, as there is a canonical 2-cocycle on the local pair groupoid (this is consistent with the fact that Fedosov’s quantization is canonically defined on a Kähler manifold, as the Levi-Civita connection is a symplectic connection). See item 8.
@@ 2 Geometric Quantization of the Lie 2-Groupoid
A Poisson structure defines a 2-cocycle in the Chevalley-Eilenberg complex of This integrates to a 2-cocycle on the 2-groupoid,121212One should interpret this as the inverse of a higher groupoid van Est map. which defines a multiplicative line bundle
@@
is given by131313We are suppressing the dependence of on in the notation.
@@ 3.3 Part 3: Representation on Sections of Prequantum Line Bundle
The construction of the prequantum line bundle is obtained by constructing the line bundle over 141414 is the pair groupoid, it has a unique arrow between any two points in as in the previous section, and pulling it back to the source fiber over We will recall the standard construction in geometric quantization below ([23]).
@@
acts by parallel transport over 151515We are implicitly using the isomorphism so that these algebroid morphisms into are the same as maps of manifolds into followed by multiplication by This descends to an action of on
@@ 4 Example: Conventional Quantum Mechanics
Consider the Kähler manifold with the standard metric.
@@
(4.0.2) which is the (signed) area of the geodesic triangle with vertices This is the natural -cocycle obtained using the Kähler structure.161616The van Est map applied to gives The multiplication of the line bundle is given by
@@
(4.0.5) -
@@
The quantization map in definition 3.2.2 is given by
@@
(4.0.9) Note that, is constant (but not covariantly constant using the canonical connection) along the polarization
(4.0.10) -
@@
The function (as defined in 4.0.7) equals the non-perturbative Moyal product ([21], [27]), which is an associative product on deforming the pointwise product and gives us lemma 1.1.1. We have that
(4.0.11) and therefore is a -algebra isomorphism between the Moyal product and its image.
@@
Explicitly (see lemma 3.1.2),
(4.0.12) This is invariant under automorphisms of the Kähler structure. The integral kernel is and is equal the integral kernel in eq. 1.1.3 (this construction doesn’t use a choice of point in ).
This completes the constructions of theorem 0.1.2. We will now discuss irreducible subrepresentations of our -algebra. -
@@
(4.0.13) for constants (not both ). Polarized sections are generated by sections of the form
for such that
(4.0.15) These give irreducible subrepresentations of the left representation of ie. is of the form 4.0.14, for any Notice that these representations come in pairs, by taking
10.
The -actions of on (as in 4.0.14) are given by191919Our conventions are such that
(4.0.16) (4.0.17) We can read off the position, momentum and Segal-Bargmann representations (ie. the holomorphic representation [13]).202020The latter is the only one which forms a -algebra subrepresentation, the former are only algebra subrepresentations, but there is essentially a unique measure turning them into -algebra representations. We are describing the representation of the quantum observables on all vector states212121A pure state is one which generates an irreducible representation, these polarized sections are pure states. However, the quantization map is defined on all sections of the prequantum line bundle, and some of these sections are mixed states. in the sense of -algebras. See [9]. before choosing a polarization. In traditional quantization, one chooses a polarization first and the Kostant-Souriau quantization prescription only quantizes a small subspace of observables which preserves polarized sections.
Generally, if then(4.0.18)
The two -representations associated to are related through integration, as in the next point.
@@
The involution lifts to an isomorphism of -representations from states of the form
(4.0.19) to states of the form
(4.0.20) The isomorphism is given simply by integration with respect to where we identify the vector spaces over and using the flat connection along the polarization, ie. under parallel transport over we have that
@@ 4.1 The Classical Limit
We can analyze the classical limit from the 2-groupoid perspective. First, note that given a Poisson structure on we get a natural Poisson structure on where the induced Poisson structure on is .222222We can replace with any subset with as an accumulation point. Then, what we have really done in this paper is integrated the extension of Lie algebroids232323 is the trivial line bundle over
@@
The case is special: isomorphism classes (ie. orbits) of objects in the 2-groupoid of are its symplectic leaves, and is the unique case where the symplectic leaves are just the points in As a result, the pure states at are just pointwise evaluation. From the 2-groupoid perspective, the measurement problem is due to there being isomorphisms between distinct points in phase space of a symplectic manifold for
@@ 5 Concluding Remarks
In our main example, one can check that if is periodic and transforms as a theta function, then transforms as a theta function. Therefore, if we carry out our construction on the symplectic torus, we will get a representation of the noncommuative torus on sections of the prequantum line bundle. It would be interesting to construct the irreducible representations.
In [22] (see also [4], [6]) it is shown that there is a sequence of “quantization maps” and representations associated to a compact Kähler manifold (Berezin-Toeplitz quantization) which induce a formal deformation quantization. However, these quantization maps are not injective — they each have finite dimensional image. We expect that our construction can be carried out in this context, and that their representations factor through our geometric quantization map with the path integrals being computed using the lattice definition in [16].

### v2

@@ Abstract
We canonically quantize a Poisson manifold to a Lie 2-groupoid, complete with a quantization map, and show that it relates geometric and deformation quantization: the perturbative expansion in of the (formal) convolution of two quantized functions yields Kontsevich’s star product. Meanwhile, we can push forward this quantization map (by integrating over homotopies of paths) to obtain a quantization map in traditional geometric quantization. This gives a polarization-free, path integral definition of the quantization map, which does not have a prior prescription for Poisson manifolds and which only has a partial prescription for symplectic manifolds. We show that in conventional quantum mechanics our operators are defined on the entire prequantum Hilbert space and that they all preserve polarized sections, contrary to the Kostant-Souriau approach.
@@ 0 Introduction
Kontsevich proved that every Poisson manifold admits a formal deformation quantization,222Obtained using a sequence of bidifferential operators. called a star product (see [15]). That is, there is an associative product on which deforms the product on in such a way that
@@
A non-perturbative approach to quantization, which was initiated by Weinstein in [26], is geometric quantization of Poisson manifolds. This quantization method aims to define a -algebra of observables — it is given by polarized sections of a twisted convolution algebra of the symplectic groupoid (the groupoid which integrates the Poisson structure, [11]). In the symplectic case, we would like this -algebra to act on polarized sections of the prequantum line bundle.333For a discussion on when a star product acts on the space of sections of such a quantization, see [18], [19].
@@
Geometric quantization is much less understood than formal deformation quantization (which is consistent with the current understanding of QFT) and is incomplete as it doesn’t prescribe a quantization map.444Here, we mean an injective -linear map from to the -algebra of observables that determines a star product. In this sense, the difficulty of geometric quantization is complementary to the difficulty of formal deformation quantization: one (partially) prescribes a -algebra but not a quantization map, the other prescribes a formal quantization map but not a -algebra.
@@
In some examples (eg. [1],[3],[22]) it has been shown that there exists a quantization map which is compatible with formal deformation quantization, but an explanation of why and when this occurs is lacking. Even in traditional geometric quantization of symplectic manifolds the quantization map is, at best, defined only on a small subspace (and assumes a polarization exists). It is desirable to have a more mathematically precise theory which directly relates these two quantization methods.
@@
The constructions we make are justified by results of the Poisson sigma model in [2], [5]. We formally define a quantization map (definition 3.2.2) by using the Lie 2-groupoid integrating a Poisson manifold ([27]) — this involves computing a path integral over homotopies of algebroid paths. In the symplectic case, this 2-groupoid is a quotient of the singular simplicial set of the manifold. Noncommutativty for arises due to distinct points in phase space being isomorphic in the 2-groupoid. This is a two dimensional sigma model explanation of quantization, somewhat as in [10].
0.1 Main Example
We compute the path integrals in the case of the pointed Kähler manifold with the canonical symplectic form (section 4).555The only examples of geometric quantization of symplectic manifolds that have been shown to have an injective quantization map are discrete quotients of and it is not difficult to compute the path integrals there. The quantization of a smooth function is a section of a trivializable line bundle over the symplectic groupoid, and can be identified with (eq. 4.0.8)
@@ 0.1 Main Example
where is the (signed) area of the polygon with vertices and the integration is over
The canonical connection on the prequantum line bundle is (definition 3.3.1) and parallel transport determines a representation of these quantized functions as integral operators on The result is
where and is the (signed) area of the geodesic triangle determined by the integral is over with respect to the product measure.
The prequantum Hilbert space contains both pure and mixed states.666In the -algebra sense. A pure vector state is one which generates an irreducible representation ([9]). Polarized sections form irreducible subrepresentations for all covariantly constant Lagrangian polarizations. That is, if is polarized (with respect to some complex linear polarization), then (eq. 4.0.19)
and this is a polarized section. In particular, we get the position, momentum and Segal-Bargmann (holomorphic) representations. The states of the latter take the form
With the Kostant–Souriau prescription, only a small subspace of functions prequantize to operators which preserve polarized sections, and this subspace depends on the polarization — only affine functions prequantize to operators which preserve all linearly polarized sections. Our operators 0.1.2 are defined polarization-free and they all preserve polarized sections, for all complex linear polarizations. These two quantizations agree on the subspace where the Kostant–Souriau prescription is defined.
In addition, with respect to the twisted convolution algebra on the groupoid and the non-perturbative Moyal product (lemma 1.1.1, eq. 4.0.11), 777The image of under the van Est map is ([17]) and is therefore independent of the complex structure.
@@
We view the role of Lagrangian polarizations in the usual definitions of geometric quantization as a tool to help guess a quantization map, but not necessarily as an essential ingredient (sometimes they don’t even exist). This is consistent with Kontsevich’s and Fedosov’s ([8]) constructions of the star product, where polarizations are absent. We will only use them to help construct irreducible representations.
In the usual approach, a common heuristic is that the prequantum Hilbert space being “too large” is a problem. However, our perspective is that the only problem is that the subalgebra of operators generated by the Kostant-Souriau prequantization map is far too large to represent observables — its image only contains first order differential operators. A precise version of this is that the Kostant-Souriau prequantization map is not compatible with a star product (definition 0.3.1). This perspective is particularly useful when the prequantum Hilbert space contains inequivalent irreducible representations.
@@ 0.2 Lie 2-Groupoid Approach to Quantization and Main Results
Its space of sections, denoted is formally a -algebra — its product is given by convolution (this involves a path integral) and its involution is complex conjugation. It’s norm is (formally) the -norm of the left regular representation.888Every algebra acts on itself on the left, forming a representation called the left regular representation. There is well-defined, canonical quantization map
@@
(formal) Let be a Poisson maniold such that the convolution of two sections of is smooth.999This should be true in a wide class of examples, including regular Poisson manifolds satisfying an integrality condition. We have that
@@
(0.2.4) of the Poisson sigma model on the disk ([5]). Therefore, perturbatively at
@@
0.3 Non-Perturbative (Tangential) Deformation Quantizations
We will discuss non-perturbative deformation quantizations. The definition we use is essentially the most general one that is compatible with both -algebraic quantization and formal deformation quantization. It sits between Rieffel’s strict quantization ([20]) and his strict deformation quantization.
@@ 0.3 Non-Perturbative (Tangential) Deformation Quantizations
be injective and - linear such that its image generates Furthermore, assume that is the inclusion map. We say that is a (non-perturbative) deformation quantization of if there is a star product on such that, for all
@@
Every regular Poisson manifold has a tangential, formal deformation quantization,101010A star product is tangential if the sequence of bidifferential operators are tangent to the symplectic leaves. see the work of Fedosov [8], also see [13], [25]. (there are a number of examples of non-regular Poisson manifolds which also have such a quantization, as discussed in [17]).111111A quantization which isn’t tangential exhibits a kind of tunneling. The maps discussed in section 0.2 have a tangential property since the orbits of the Lie 2-groupoid are the symplectic leaves.
@@
where the supremum is over all symplectic leaves and is the -norm of pulled back to 121212With respect to the measure induced by the symplectic form. We denote this space by
@@
This would give a a non-perturbative form of Kontsevich’s star product on such Poisson manifolds.131313For examples of non-tangential quantizations, see the quantizations of the dual of nilpotent Lie algebras [21].
@@
In formal deformation quantization, it’s really the gauge equivalence class of the formal quantization map that’s important. Similarly, we might define two quantization maps to be gauge equivalent if there is an automorphism of -algebras intertwining them.
@@ 1 The Star Product and the Poisson Sigma Model
The star product is then given by141414More precisely, we can say that the perturbative expansions at is Kontsevich’s star product.
@@ 1.1 The Tangential Star Product
A lattice approach to defining these path integrals is described in [17], see appendix 5. On a Kähler manifold this kernel should be canonically defined, as there is a canonical 2-cocycle on the local pair groupoid (this is consistent with the fact that Fedosov’s quantization is canonically defined on a Kähler manifold, as the Levi-Civita connection is a symplectic connection). See LABEL:moyal.
@@ 2 Geometric Quantization of the Lie 2-Groupoid
A Poisson structure defines a 2-cocycle in the Chevalley-Eilenberg complex of This integrates to a 2-cocycle on the 2-groupoid,151515One should interpret this as the inverse of a higher groupoid van Est map. which defines a multiplicative line bundle
@@
is given by161616We are suppressing the dependence of on in the notation.
@@ 3.3 Part 3: Representation on Sections of Prequantum Line Bundle
The construction of the prequantum line bundle is obtained by constructing the line bundle over 171717 is the pair groupoid, it has a unique arrow between any two points in as in the previous section, and pulling it back to the source fiber over We will recall the standard construction in geometric quantization below ([24]).
@@
acts by parallel transport over 181818We are implicitly using the isomorphism so that these algebroid morphisms into are the same as maps of manifolds into followed by multiplication by This descends to an action of on
@@ 4 Example: Conventional Quantum Mechanics
Consider the pointed Kähler manifold 191919All of our constructions will depend only on this pointed Kähler structure.
@@
(4.0.2) which is the (signed) area of the geodesic triangle with vertices This is the natural -cocycle obtained using the Kähler structure.202020The van Est map applied to gives The multiplication of the line bundle is given by
@@
(4.0.5) We have implicitly used the “geodesic” trivialization of in step 2.
@@
The quantization map in definition 3.2.2 is given by212121We have normalized by dividing out by eq. 1.1.3.
@@
(4.0.9) Thus, equivariant under automorphisms (of the pointed Kähler structure).
@@
Note that there is a Lagrangian polarization induced by the morphism
(4.0.10) and is polarized with respect to the connection 222222Compare this with the construction of the quantization in [11], for which a polarization needs to be chosen. The two quantization maps are related by an automorphism of
@@
The product (defined in lemma 1.1.1) is given by
(4.0.11) (4.0.12) This is equal to in theorem 0.2.2 and is the non-perturbative Moyal product ([22], [28]). It is invariant under automorphisms of the Kähler manifold and the integral kernel is which is equal the integral kernel in eq. 1.1.3.
@@
We have that and in fact
This completes the constructions of theorem 0.2.2. We will now discuss irreducible subrepresentations of our -algebra, using eq. 4.0.13.
10.
@@
(4.0.14) for constants (not both ). A polarized section can be written as
for any such that
(4.0.16) With respect to the Kähler polarization, polarized sections are of the form
These give irreducible subrepresentations of the left representation of ie. is of the form 4.0.15, for the same and for any See eq. 4.0.19. -
@@
The actions of the quantizations of on (as in 4.0.15) are given by242424Our conventions imply that
(4.0.17) (4.0.18) We can read off the position, momentum and Segal-Bargmann representations (ie. the holomorphic representation [14]).252525The holomorphic representation is the only one which forms a -algebra subrepresentation, the former are only algebra subrepresentations, but there is essentially a unique measure turning them into -algebra representations.
We are describing the representation of the quantum observables on all vector states262626A pure state is one which generates an irreducible representation, these polarized sections are pure states. However, the quantization map is defined on all sections of the prequantum line bundle, and some of these sections are mixed states. in the sense of -algebras. See [9]. before choosing a polarization. In traditional quantization, one chooses a polarization first and the Kostant-Souriau quantization prescription only quantizes a small subspace of observables which preserves polarized sections.
Generally, if is a polarized section, then(4.0.19) -
12.
The involution lifts to an isomorphism of representations from states polarized with respect to to states polarized with respect to The isomorphism is given simply by integration with respect to where we identify the vector spaces over and using the flat connection along the polarization, ie. under parallel transport over we have that
@@ 4.1 The Classical Limit
We can analyze the classical limit from the 2-groupoid perspective. First, note that given a Poisson structure on we get a natural Poisson structure on where the induced Poisson structure on is .272727We can replace with any subset with as an accumulation point. Then, what we have really done in this paper is integrated the extension of Lie algebroids282828 is the trivial line bundle over
@@
The case is special: isomorphism classes (ie. orbits) of objects in the 2-groupoid of are its symplectic leaves, and is the unique case where the symplectic leaves are just the points in As a result, the pure states at are just pointwise evaluation. From the 2-groupoid perspective, the measurement problem is due to the existence of isomorphisms between distinct points in phase space of a symplectic manifold, for
@@ 5 Concluding Remarks
In our main example, one can check that if is periodic and transforms as a theta function under translations, then transforms as a theta function. Therefore, if we carry out our construction on the symplectic torus, we will get a representation of the noncommuative torus on sections of the prequantum line bundle. It would be interesting to construct the irreducible representations.
In [23] (see also [4], [6]) it is shown that there is a sequence of “quantization maps” and representations associated to a compact Kähler manifold (Berezin-Toeplitz quantization) which induce a formal deformation quantization. However, these quantization maps are far from injective. We expect that our construction can be carried out in this context, and that their representations factor through our geometric quantization map with the path integrals being computed as in [17].
@@
The path integrals in 4 can then be computed using the integral representation of the delta function.